OpenStudy (anonymous):

1. the probability that a dessert sold at a certain chocolate is 74% the probability that a dessert contians chocolate and nuts is 24% find the probability that a randomly chosen chocolate dessert contains nuts round to the nearest tenth of a percent. A. 17.8% B. 32.4% C. 31.4% D. 50%

OpenStudy (atlas):

the probability of finding a dessert with chocolates AND nuts was 24%.......which means 24 out of a 100 and you know that the chosen dessert is chocolate - so you have to find the probabilty of it having nut - so you have to convert the 24out of 100 chance to x out of 74 24/100 = x/74 or, x = 0.24 * 74 = 17.8